OpenStudy (anonymous):

Which is an equation of the circle with a radius of 9 units and center at (–4, 2)?

OpenStudy (anonymous):

A. (x − 9)2 + (y + 4)2 = 4 B. (x − 4)2 + (y + 2)2 = 81 C. (x + 4)2 + (y − 2)2 = 81 D. (x − 2)2 + (y + 4)2 = 81

Directrix (directrix):

Write your circle as shown in the attachment.

Directrix (directrix):

Post what you get.

OpenStudy (anonymous):

(x + 4)2 + (y − 2)2 = 81

Directrix (directrix):

That is what I got, too.